Name: Date: CHE 112L Pre-Lab Assignment Experiment # 2: Equilibrium Constant for a Chemical Reaction Assignment is due at the beginning of next lab session Constant t lab session SHOW WORK. 1. Write the equilibrium constant expression for the reaction below (Kay) CH4(g) + 2 H2S(g) ←→ CS2(g) + 4 H2(g) A reaction mixture initially contains 0.50 M CH4 and 0.75 M H2S. If the equilibrium concentration of H2 is 0.44 M, find the equilibrium constant (Ke) for the reaction. 2. Express the equilibrium constant for the following reaction. P4(s) +502(g) ←→ P4010(s)
Answer #1

1.

                              CH4 + 2H2S <--> CS2 +4H2
initial:                      0.5      0.75           0       0
at equilibrium:       0.5-x     0.75-2x        x      4x

Given that equilibrium concentration of H2= 0.44 M
so, 4x = 0.44 M
x = 0.44/4 = 0.11 M

At equilibrium concentration of various species are:
[CH4] = 0.5-x = 0.5 - 0.11 = 0.39 M
[H2S] = 0.75 -2x = 0.75 - 2*0.11 = 0.53 M
[CS2] = x = 0.11 M
[H2]= 4x = 4*0.11 =0.44 M


Kc = [H2]^4 * [CS2] / [CH4] [H2S]^2
   = (0.44)^4 *(0.11) /(0.39)*(0.53)^2
   = 0.0376

Ans: 0.0376

2.

Equilibrium constant expression of Kc will be:

Kc= [P4O10] / {[P4]*[O2]^5}

Sinceconcentration of solid doesn't change, we can write Kc as:

Kc= 1/ [O2]5

   = [O2]-5